# Deep Common Feature Mining for Efficient Video Semantic Segmentation

### v1

@@ Abstract
Recent advancements in video semantic segmentation have made substantial progress by exploiting temporal correlations. Nevertheless, persistent challenges, including redundant computation and the reliability of the feature propagation process, underscore the need for further innovation. In response, we present Deep Common Feature Mining (DCFM), a novel approach strategically designed to address these challenges by leveraging the concept of feature sharing. DCFM explicitly decomposes features into two complementary components. The common representation extracted from a key-frame furnishes essential high-level information to neighboring non-key frames, allowing for direct re-utilization without feature propagation. Simultaneously, the independent feature, derived from each video frame, captures rapidly changing information, providing frame-specific clues crucial for segmentation. To achieve such decomposition, we employ a symmetric training strategy tailored for sparsely annotated data, empowering the backbone to learn a robust high-level representation enriched with common information. Additionally, we incorporate a self-supervised loss function to reinforce intra-class feature similarity and enhance temporal consistency. Experimental evaluations on the VSPW and Cityscapes datasets demonstrate the effectiveness of our method, showing a superior balance between accuracy and efficiency.
Figure 1: Trade-off illustration: various semantic segmentation methods using MiT backbones on VSPW validation set [28], including SegFormer (image baseline) [43], CFFM [36], MRCFA [37], and DCFM (Ours). By adjusting the keyframe interval $K$ during inference, DCFM achieves impressive speed (80 FPS @ $K$=10) while maintaining high level of accuracy (45.4% mIoU).
1 Introduction
@@
Semantic segmentation is a prominent dense prediction task in the field of computer vision, involving the assignment of pixel-wise category labels to images. Significant progress in this field has been propelled by the exploration of robust deep neural backbones [32, 54, 14, 40, 43, 9], elaborate context aggregation strategies [5, 6, 7, 50, 47, 24, 43, 9], and additional boundary supervision [5, 33, 48]. In recent years, the proliferation of real-world applications, ranging from autonomous driving to robot sensing, has intensified the focus on video semantic segmentation (VSS). The success of this task extends beyond achieving high accuracy [13, 30, 36, 37, 2], necessitating a trade-off between inference efficiency [46, 12, 31, 55, 45, 15, 1] and temporal consistency [25, 49].
@@
While it is straightforward to extend the capabilities of contemporary image models to VSS, the naive adaptation to every video frame introduces considerable computational overhead and neglects the inherent temporal correlations in videos, which are crucial for understanding scenes with lower uncertainty and enhancing temporal smoothness [25, 49]. To address this, methods like LMANet [30], CFFM [36], and MRCFA [37] explicitly aggregate information over multiple frames, utilizing Transformers [38] or spatial-temporal attention for context encoding. Despite their effectiveness, the heavy computational cost and complex temporal modeling hinder practical deployment [13, 39, 19, 30, 36, 37]. As illustrated in Figure 1, achieving high accuracy with CFFM and MRCFA in certain experimental environment leads to a slow inference speed of approximately 6 FPS.
@@
On the flip side, the temporal continuity in videos presents an opportunity to enhance efficiency. Acknowledging the frequent redundancy in content across adjacent frames, a promising approach is to reuse high-level features extracted from selected keyframes, striking a refined balance between inference speed and accuracy [31, 55, 34, 45, 1, 16]. This can be achieved by leveraging a pre-trained optical flow model for feature warping [17] or a well-designed propagation module [31, 1, 16] to transmit features or segmentation maps from keyframes to adjacent ones, significantly reducing redundant computation. However, existing keyframe-based methods tend to excessively rely on the propagation process, leaving accurate feature prediction unsolved, especially in complex scenarios involving large motions and severe occlusions [31, 52, 49]. Furthermore, propagating unstable features tends to accumulate errors [44].
@@
In this study, our emphasis is on achieving efficient VSS by learning a discriminative representation to assist adjacent frames in perceiving the scene without resorting to time-consuming re-calibration operations like optical-flow-based warping. To address both aspects, we propose decomposing the entangled feature into two complementary components: (1) the common representation, providing essential information for segmentation and directly reusable across frames without feature propagation. This corresponds primarily to high-level semantics that remain relatively constant over time; and (2) the independent representation, easily extractable to offer frame-specific clues for accurate segmentation. Figure LABEL:fig:moti visualizes the distinctions between our approach and the existing ones.
@@
2 Related work
2.1 Image semantic segmentation
Fully Convolutional Networks (FCN) [27] paved the way for the application of Convolutional Neural Networks (CNNs) in image semantic segmentation. Recognizing the necessity for informative spatial contexts, subsequent methods focused on refining network architectures to aggregate rich contextual information. Among them, the encoding-decoding structures have proven to be particularly effective [32, 54, 8, 3]. PSPNet [50], along with studies like DeepLabV3 [7] and OCRNet [47], introduced post-processing modules for enhanced spatial awareness. HRNet-V2 [40] utilized exchange blocks for multi-scale information fusion. The application of vision transformers in image tasks, as seen in SETR [53] and Segformer [43], further advanced this field. Additionally, MaskFormer and Mask2Former improved DETR [4] for segmentation tasks, steering away from pixel-by-pixel classification. Some methods also incorporated boundary supervision [5, 33, 48] to enhance accuracy.
Another notable trend aimed at designing efficient networks for practical deployment [46, 21, 22, 51]. In particular, STDC [12] removed structural redundancy from BiSeNet [46] and introduced a new backbone, considering features extracted by shallow layers as the output of the spatial path, thereby further reducing model complexity.
2.2 Video semantic segmentation
To exploit temporal correlations in videos, various methods [13, 18, 19, 29, 15, 39, 30, 36, 37, 2, 23] explicitly aggregate high-level features across frames for improved accuracy. Attention mechanisms [38] and vision transformers [11], recognized for their proficiency in long-term dependency modeling, are prominent in VSS. TDNet [15] and TMANet [39], for instance, use attention-based approaches for feature aggregation, while CFFM [36] employs a transformer to learn a unified representation of static and motion contexts. Despite their accuracy, these methods often demand substantial computational resources.
Another category of methods explores the advantages of video continuity, prioritizing a favorable speed-accuracy trade-off [31, 55, 34, 45, 41, 16]. Typically, a robust baseline segmentation model is selectively applied to keyframes, and the extracted high-level features are reused by subsequent non-key frames, effectively reducing average latency. However, an additional propagation module is generally required, introducing extra computational overhead.
@@ 2.2 Video semantic segmentation
3 Method
This section describes the DCFM approach, delving into its network architecture, efficient inference pipeline, and the innovative training strategy with a specialized loss function.
3.1 Overview
We employ a contemporary image segmentation model as the backbone. As observed in Clock [34] and STDC [12], features of the deep layers of the backbone tend to capture high-level semantics, evolving gradually over time. In contrast, shallow features excel at depicting details with quicker update rates. During training, we emphasize this property through common feature mining, ensuring deep representations from keyframes can be effortlessly reused across frames during inference, eliminating the need for re-calibration. This achieves an optimal balance between inference speed and accuracy. To accomplish this, we group the deep and shallow layers of the backbone network into two stages to capture the common and independent representations, respectively. The symmetric training strategy alternates between a labeled keyframe and a non-key frame, facilitating high-level common feature mining (detailed in Section 3.2). Additionally, a proposed self-supervised loss minimizes perceptual distance for intra-class pixels, enhancing representation robustness (detailed in Section 3.3).
In Figure LABEL:fig:framework, we present the complete network architecture when the model is applied to a video. For a video comprising frames $\{x_{0},x_{1},...,x_{i},...\}$, where $x_{i}\in\mathbb{R}^{C\times H\times W}$, we select one keyframe for every $K$ frames. The shallow and deep layers of the backbone are organized to extract features from this keyframe $x_{i}$, expressed as:
@@ 3.1 Overview
[…]
@@
Feature fusion module. The structure of FFM is illustrated in Figure LABEL:fig:ffm. Initially, we align the sizes of common and independent feature maps through interpolation. To alleviate computational burdens and prevent short-distance back-propagation from affecting training, we concatenate $F^{i}_{co}$ with half of the independent feature $F^{i}_{in}$ (or $F^{j}_{in}$). Then we employ only one convolution layer for feature fusion, ensuring minimal computational load.
3.2 Deep common feature mining
@@
3.3 Self-supervised consistency loss
@@
3.4 Adaptive keyframe scheduling
Previous research [31, 45, 55] has made valuable contributions to adaptive keyframe scheduling (AKS), allowing keyframe-based VSS methods to adapt to dynamic scenes with varying motion levels instead of using a fixed keyframe scheduling. While DCFM does not explicitly integrate a keyframe scheduling module, we use a simple AKS scheme [34] to enhance flexibility. We calculate a score between each frame and the previous keyframe during inference, defined as $s_{j}=\text{mean}(|x_{i}-x_{j}|)$, where $x_{i}$ is the previous keyframe, and $j>i$. We traverse through $x_{j}$ until the score $s_{j}$ exceeds a predefined threshold $S$ or reaches the end of the video. This adaptive keyframe selection strategy allows DCFM to adapt to varying motion levels, providing flexibility in handling dynamic scenes.
Methods Backbone mIoU $\uparrow$ wIoU $\uparrow$ $\text{mVC}_{8}$uparrow $\text{mVC}_{16}$uparrow Params $\downarrow$ GFLOPS $\downarrow$ FPS $\uparrow$
@@ 3.4 Adaptive keyframe scheduling
DCFM (Ours, $K$=2) MiT-B2 43.7 63.7 87.7 83.2 24.8M 22.9 67.6
DCFM (Ours, $K$=6) MiT-B2 42.4 62.6 90.4 86.2 24.8M 12.3 113.6
@@
DCFM (Ours, $K$=2) MiT-B5 48.2 65.5 89.0 85.0 82.1M 57.0 25.4
DCFM (Ours, $K$=6) MiT-B5 46.5 64.1 91.7 88.0 82.1M 23.6 58.8
Table 1: Comparison of accuracy (mIoU and wIoU), temporal consistency (mVC), parameter number, GFLOPs, and inference speed (FPS) on the VSPW [28] validation dataset. Figure LABEL:fig:sota-vs provides a visual illustration, focusing on two dimensions: mIoU and FPS. $K$ indicates the keyframe interval during inference.
4 Experiments
4.1 Datasets and evaluation metrics

### v2

@@ Abstract
Recent advancements in video semantic segmentation have made substantial progress by exploiting temporal correlations. Nevertheless, persistent challenges, including redundant computation and the reliability of the feature propagation process, underscore the need for further innovation. In response, we present Deep Common Feature Mining (DCFM), a novel approach strategically designed to address these challenges by leveraging the concept of feature sharing. DCFM explicitly decomposes features into two complementary components. The common representation extracted from a key-frame furnishes essential high-level information to neighboring non-key frames, allowing for direct re-utilization without feature propagation. Simultaneously, the independent feature, derived from each video frame, captures rapidly changing information, providing frame-specific clues crucial for segmentation. To achieve such decomposition, we employ a symmetric training strategy tailored for sparsely annotated data, empowering the backbone to learn a robust high-level representation enriched with common information. Additionally, we incorporate a self-supervised loss function to reinforce intra-class feature similarity and enhance temporal consistency. Experimental evaluations on the VSPW and Cityscapes datasets demonstrate the effectiveness of our method, showing a superior balance between accuracy and efficiency. The implementation is available at https://github.com/BUAAHugeGun/DCFM.
Index Terms:
Video semantic segmentation, efficient segmentation, keyframe-based model, common feature mining, temporal consistency.
I Introduction
@@
Semantic segmentation is a prominent dense prediction task in the field of computer vision, involving the assignment of pixel-wise category labels to images. Substantial progress in this field has been propelled by the exploration of robust deep neural backbones [1, 2], elaborate context aggregation strategies [3, 4], and additional boundary supervision [5, 6]. In recent years, real-world applications like autonomous driving and robot sensing have heightened interest in video semantic segmentation (VSS). Success in VSS goes beyond high accuracy  [7, 8], necessitating a trade-off between accuracy, inference efficiency [9, 10] and temporal consistency [12, 13].
Figure 1: The trade-off between speed and accuracy is illustrated through various semantic segmentation methods using MiT backbones on the VSPW validation set [14]. These methods include SegFormer (image baseline) [2], CFFM [15], MRCFA [16], and our proposed method, DCFM. By adjusting the keyframe interval $K$ during inference, DCFM achieves impressive speed (80 FPS @ $K$=10) while maintaining a high level of accuracy (46% mIoU).
(a) Per-frame inference
(b) Feature aggregation
(c) Feature propagation
(d) DCFM (Ours)
Figure 2: Video semantic segmentation pipelines. (a) Applying an image segmentation model to each frame independently; (b) Aggregating features from multiple frames to predict the segmentation map for a target frame, enhancing accuracy; (c) Holistically propagating the feature of a keyframe to subsequent non-key frames for reuse, improving efficiency; and (d) The proposed method decomposes the feature into two complementary parts, enabling direct reuse of the common part by non-key frames without re-calibration, improving efficiency.
@@
While it is straightforward to extend the capabilities of contemporary image models to VSS, the naive adaptation to every video frame introduces considerable computational overhead and neglects the inherent temporal correlations in videos, which are crucial for understanding scenes with lower uncertainty and enhancing temporal smoothness [12, 13]. To address this, methods like CFFM [15], MRCFA [16], and LMANet [17] explicitly aggregate information over multiple frames, utilizing Transformers [18] or spatial-temporal attention for context encoding. Despite their effectiveness, the heavy computational cost and complex temporal modeling hinder practical deployment [7, 17]. As illustrated in Figure 1, achieving high accuracy with CFFM and MRCFA in a certain experimental environment leads to a slow inference speed of approximately 6 FPS.
@@
On the flip side, the temporal continuity in videos presents an opportunity to enhance efficiency. Acknowledging the frequent redundancy in content across adjacent frames, a promising approach is to reuse high-level features extracted from selected keyframes, striking a refined balance between inference speed and accuracy [19, 20]. This can be achieved by leveraging a pre-trained optical flow model for feature warping [21, 22, 23] or a well-designed propagation module [9, 10, 11] to transmit features or segmentation maps from keyframes to adjacent ones, significantly reducing redundant computation. However, existing keyframe-based methods tend to excessively rely on the propagation process, leaving accurate feature prediction unsolved, especially in complex scenarios involving large motions and severe occlusions [9, 24, 13]. Furthermore, propagating unstable features tends to accumulate errors [25].
@@
In this study, our emphasis is on achieving efficient VSS by learning a discriminative representation to assist adjacent frames in perceiving the scene without resorting to time-consuming re-calibration operations like optical-flow-based warping. To address both aspects, we propose decomposing the entangled feature into two complementary components: (1) the common representation, providing essential information for segmentation and directly reusable across frames without feature propagation. This corresponds primarily to high-level semantics that remain relatively constant over time; and (2) the independent representation, easily extractable to offer frame-specific clues for accurate segmentation. Figure 2 visualizes the distinctions between our approach and the existing ones.
@@
II Related work
II-A Image semantic segmentation
Fully Convolutional Networks (FCN) [27] paved the way for the application of Convolutional Neural Networks (CNNs) in image semantic segmentation by replacing fully connected layers with convolution layers. Recognizing the necessity for informative spatial contexts, subsequent methods focused on refining network architectures to aggregate rich contextual information. Among them, the encoding-decoding structures have proven to be particularly effective [28, 29, 30, 31, 32]. PSPNet [33], along with studies like DeepLabV3 [34] and OCRNet [35], introduced post-processing modules for enhanced spatial awareness. HRNet-V2 [36] utilized exchange blocks for multi-scale information fusion, while PSA [37] explored attention mechanisms within backbone blocks. The application of vision transformers in image tasks, as seen in Structtoken [38], SETR [39], Graph-seg [40] and Segformer [2], further advanced this field. Additionally, MaskFormer [41] and Mask2Former [42] improved DETR [43] for segmentation tasks, steering away from pixel-by-pixel classification. Some methods also incorporated boundary supervision [44, 5, 6, 32] to enhance accuracy. In addition, a number of methods [45, 46, 47, 48, 49] employ domain adaptation techniques to utilize additional semantic segmentation datasets, thereby aiding segmentation on the target domain.
Another notable trend aimed at designing efficient networks for practical deployment [50, 51, 52, 53, 54]. In particular, BiSeNet [50] aggregated feature maps from spatial and context paths to improve accuracy in lightweight networks. STDC [55] removed structural redundancy from BiSeNet and introduced a new backbone, considering features extracted by shallow layers as the output of the spatial path, thereby further reducing model complexity.
II-B Video semantic segmentation
To exploit temporal correlations in videos, various methods [7, 56, 57, 58, 59, 60, 17, 15, 16, 8, 61] explicitly aggregate high-level features across frames for improved accuracy. Attention mechanisms [18] and vision transformers [62], recognized for their proficiency in long-term dependency modeling, are prominent in VSS. TDNet [59] and TMANet [60], for instance, use attention-based approaches for feature aggregation, while CFFM [15] employs a transformer to learn a unified representation of static and motion contexts.
Another category of methods explores the advantages of video continuity, prioritizing a favorable speed-accuracy trade-off [9, 63, 19, 20, 64, 11, 65]. Typically, a robust baseline segmentation model is selectively applied to keyframes, and the extracted high-level features are reused by subsequent non-key frames, effectively reducing average latency.
@@ II-B Video semantic segmentation
(a) Framework Overview
(b) Feature Fusion Module
Figure 3: Overview of Deep Common Feature Mining (DCFM) approach. (a) The complete network architecture for video segmentation during inference. The backbone, based on an image semantic segmentation network, groups layers into two stages to output common and independent features, respectively. The extracted common representation $F_{co}$ can be explicitly reused across frames for high efficiency. (b) Structure of the Feature Fusion Module (FFM), where the exemplar target frame is a non-key frame.
III Method
This section elaborates on the DCFM approach, detailing its network architecture, efficient inference pipeline, and the innovative training strategy incorporating a specialized loss function.
III-A Overview
We employ a contemporary image segmentation model as the backbone. As observed in Clock [19] and STDC [55], features of the deep layers of the backbone tend to capture high-level semantics, evolving gradually over time. In contrast, shallow features excel at depicting details with quicker update rates. Therefore, we divide the model into two parts: $\text{Enc}(x)=\text{Enc}_{hi}(\text{Enc}_{lo}(x))$. During training, we emphasize this property through common feature mining, ensuring deep representations from keyframes can be effortlessly reused across frames during inference, eliminating the need for re-calibration. This achieves an optimal balance between inference speed and accuracy. To accomplish this, we group the deep and shallow layers of the backbone network into two stages to capture the common and independent representations, respectively. The symmetric training strategy alternates between a labeled keyframe and a non-key frame, facilitating high-level common feature mining (detailed in Section III-B). Additionally, a proposed self-supervised loss minimizes perceptual distance for intra-class pixels, enhancing representation robustness (detailed in Section III-C).
In Figure 3a, we present the complete network architecture when the model is applied to a video. For a video comprising frames $\{x_{0},x_{1},...,x_{i},...\}$, where $x_{i}\in\mathbb{R}^{C\times H\times W}$, we select one keyframe for every $K$ frames. The shallow and deep layers of the backbone are organized to extract features from this keyframe $x_{i}$, expressed as:
@@ III-A Overview
where $F^{i}_{co}\in\mathbb{R}^{C_{F}\times H_{F}\times W_{F}}$ and $F^{i}_{in}\in\mathbb{R}^{C_{f}\times H_{f}\times W_{f}}$ denote the deep common feature and the independent feature extracted from two certain layers of the network, respectively. Correspondingly, Enc and $\textit{Enc}_{lo}$ represent the complete encoder and its early stage. Conv indicates a convolution layer, mapping the encoded deep features to a fixed dimension $C_{F}$. The two types of spatially misaligned features are combined using a lightweight feature fusion module, denoted as FFM, to create a more comprehensive representation, followed by a segmentation head, also known as the decoder, to produce the result $\hat{y}^{i}_{I}$:
@@
Feature fusion module. The structure of FFM is illustrated in Figure 3b. Initially, we align the sizes of common and independent feature maps through interpolation. To alleviate computational burdens and prevent short-distance back-propagation from affecting training, we concatenate $F^{i}_{co}$ with half of the independent feature $F^{i}_{in}$ (or $F^{j}_{in}$). Then we employ only one convolution layer for feature fusion, ensuring minimal computational load.
III-B Deep common feature mining
@@
III-C Self-supervised consistency loss
@@
III-D Adaptive keyframe scheduling
Previous research [9, 20, 63] has made valuable contributions to adaptive keyframe scheduling (AKS), allowing keyframe-based VSS methods to adapt to dynamic scenes with varying motion levels instead of using a fixed keyframe scheduling. While DCFM does not explicitly integrate a keyframe scheduling module, we use a simple AKS scheme [19] to enhance flexibility. We calculate a score between each frame and the previous keyframe during inference, defined as $s_{j}=\text{mean}(|x_{i}-x_{j}|)$, where $x_{i}$ is the previous keyframe, and $j>i$. We traverse through $x_{j}$ until the score $s_{j}$ exceeds a predefined threshold $S$ or reaches the end of the video. This adaptive keyframe selection strategy allows DCFM to adapt to varying motion levels, providing flexibility in handling dynamic scenes. Furthermore, to facilitate comparison with experiments utilizing a fixed keyframe interval, we introduce the concept of the minimum keyframe interval $\text{Min}_{K}$. We then compare various values of $\text{Min}_{K}$ with models employing fixed keyframe intervals in our experiments. Algorithm 1 details the inference process of DCFM, where $t_{p}$ and $t_{s}$ denote the indices of the nearest keyframes in the previous and future frames, respectively. Correspondingly, $F_{p}$ and $F_{s}$ represent the deep features of these two keyframes, and $f$ denote the shallow feature of the current frame. As shown in the algorithm, a simple AKS strategy can be easily integrated into the DCFM inference process.
Input: Video frames $\{x_{1},x_{2},...,x_{n}\}$; Encoder $\text{Enc}=\text{Enc}_{hi}\circ\text{Enc}_{lo}$; Decoder Dec; Minimum keyframe interval $\text{Min}_{K}$; AKS threshold $S$; The index of the first keyframe $T$; Feature fusion module FFM;
Output: Segmentation results: $\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{n}\}$;
1 $t_{p}\leftarrow 0$, $t_{s}\leftarrow T$;
[…]
5for $i\leftarrow 1$ to $n-1$ do
6       if $i=t_{s}$ then
7             $t_{p}\leftarrow t_{s}$;
8            $F_{p}\leftarrow F_{s}$;
9            $t_{s}\leftarrow t_{s}+\text{Min}_{K}$;
10            while $t_{s}<n$ and $\text{mean}(|x_{t_{s}}-x_{t_{p}}|)\leq S$ do
11                   $t_{s}\leftarrow t_{s}+1$;
12             end while
13
14            $F_{s}\leftarrow\text{Norm}(\text{Conv}(\text{Enc}(x_{t_{s}}))$;
15            $f\leftarrow\text{Norm}(\text{Enc}_{lo}(x_{t_{s}}))$;
16            $\hat{y}_{t_{s}}\leftarrow\text{Dec}(\text{FFM}(F_{s},f))$;
17       end if
18      else
19             $f\leftarrow\text{Norm}(\text{Enc}_{lo}(x_{i}))$;
20            $\hat{y}_{i}\leftarrow\frac{1}{2}(\text{Dec}(\text{FFM}(F_{p},f))+\text{Dec}(\text{FFM}(F_{s},f)))$;
21       end if
22
23 end for
Algorithm 1 Inference process of DCFM with a simple Adaptive Keyframe Schedulin (AKS).
TABLE I: Detailed experimental setup, where $C_{out}$ indicates the number of original feature channels produced by the backbone network.
Dataset VSPW [14] Cityscapes [26] CamVid [69]
Backbone MiT [2] MiT [2] ResNet [1] ResNet [1]
Crop size $480\times 853$ $512\times 1024$ $768\times 1536$ $640\times 640$
Test size $480\times 853$ $512\times 1024$ $768\times 1536$ $720\times 960$
Batch size 16 16 8 8
Optimizer AdamW AdamW SGD SGD
$C_{F}$ $C_{out}$ $C_{out}$ $\frac{5}{16}C_{out}$ $\frac{5}{16}C_{out}$
$\lambda_{B}$ 0.2 0.4 0.4 0.4
$\lambda_{C}$ 5.0 10.0 10.0 10.0
LR (Enc) 3e-5 6e-5 1e-2 2e-2
LR (Dec) 3e-4 6e-4 1e-2 2e-2
Scheduler Poly (P=1) Poly (P=1) Poly (P=0.9) Poly (P=0.9)
Iterations 160k 160k 80k 40k
Warm up $\times$
Post-proc. $\times$ $\times$ PPM [33] PPM [33]
$\text{Enc}_{lo}$ $1^{st}$ ViT block $1^{st}$ ViT block layer1 layer1
TABLE II: Comparison of accuracy (mIoU and wIoU), temporal consistency (mVC), parameter number, GFLOPs, and inference speed (FPS) on the VSPW [14] validation dataset. Figure 6a provides a visual illustration, focusing on two dimensions: mIoU and FPS. $K$ indicates the keyframe interval during inference. We measure the computational costs in FLOPs (G), averaged over videos with a resolution of 480 × 853 pixels.
Methods Backbone mIoU(%) $\uparrow$ wIoU(%) $\uparrow$ $\text{mVC}_{8}$(%) $\uparrow$ $\text{mVC}_{16}$(%) $\uparrow$ Params $\downarrow$ GFLOPs $\downarrow$ FPS $\uparrow$
@@ 20            $\hat{y}_{i}\leftarrow\frac{1}{2}(\text{Dec}(\text{FFM}(F_{p},f))+\text{Dec}(\text{FFM}(F_{s},f)))$;
DCFM (Ours, $K$=1) MiT-B2 45.0 64.3 86.6 82.0 24.8M 38.9 42.0
DCFM (Ours, $K$=2) MiT-B2 44.8 64.2 88.1 83.5 24.8M 22.9 67.6
DCFM (Ours, $K$=6) MiT-B2 43.5 63.0 90.8 86.5 24.8M 12.3 113.6
@@
DCFM (Ours, $K$=1) MiT-B5 49.5 66.0 89.0 85.5 82.1M 107.0 13.7
DCFM (Ours, $K$=2) MiT-B5 49.3 65.9 90.2 86.7 82.1M 57.0 25.4
DCFM (Ours, $K$=6) MiT-B5 47.6 64.7 92.3 88.9 82.1M 23.6 58.8
IV Experiments
DCFM is agnostic to specific models and seamlessly integrates with SOTA backbones. We illustrate its versatility by applying it to MiT (Transformer) and ResNet (CNNs). We conduct comprehensive evaluations on two benchmarks, namely VSPW [14] and Cityscapes [26]. Our model is compared with state-of-the-art methods, and we include ablation studies to validate the specifically designed network architecture and training loss.
IV-A Datasets and evaluation metrics